OpenStudy (meilendurcer):

What is 0.26 written as a fraction in simplest form?

jimthompson5910 (jim_thompson5910):

0.26 = 26/100 = ??? I'll let you reduce that completely

OpenStudy (meilendurcer):

13/25?

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

what factors do 26 and 100 have in common?

OpenStudy (meilendurcer):

13/50???

jimthompson5910 (jim_thompson5910):

you got it, 0.26 = 26/100 = 13/50

OpenStudy (meilendurcer):

thanks

jimthompson5910 (jim_thompson5910):

np

OpenStudy (meilendurcer):

another one? What is \[\frac{ 5 }{ 9 }\] written as a decimal?

OpenStudy (mathstudent55):

Divide it using long division, or use a calculator.